### v2

@@ 1 Introduction
Since a signature-type changing metric is necessarily either degenerate or discontinuous at the locus of signature change [9], we will allow for the metric to become degenerate. Hence, in the present article we will discuss singular semi-Riemannian manifolds for which the metric constitutes a smooth (0,2)(0,2)–tensor field that is degenerate at a subset $\mathcal{H}\subset M$, where the bilinear type of the metric changes upon crossing $\mathcal{H}$.
@@ 1.1 Transverse type-changing singular semi-Riemannian manifold
We align with [25] in requiring that $(M,g)$ be a semi-Riemannian manifold with $\dim M\geq 2$, where $g$ is a smooth, symmetric, degenerate (0,2)(0,2)–tensor on $M$, and $\mathcal{H}:=\{q\in M\!\!:g\!\!\mid_{q}is\;degenerate\}$. This means $\mathcal{H}$ is the locus where the rank of $g$ fails to be maximal. In addition, we assume that one connected component of $M\setminus\mathcal{H}$ is Riemannian, denoted by $M_{R}$, while all other connected components $(M_{L_{\alpha}})_{\alpha\in I}\subseteq M_{L}\subset M$ are Lorentzian, where $M_{L}:=\underset{\alpha\in I}{\bigcup}M_{L_{\alpha}}$represents the Lorentzian domain. Furthermore, we assume throughout that the point set $\mathcal{H}$, where $g$ becomes degenerate is not empty.
@@
    We call the metric $g$ a codimension–$1$ transverse type-changing metric if $d(\det([g_{\mu\nu}]))_{q}\neq 0$ for any $q\in\mathcal{H}$ and any local coordinate system $\xi=(x^{0},\ldots,x^{n-1})$ around $q$. Then we call $(M,g)$ a transverse type-changing singular semi-Riemannian manifold [1, 25]
@@
Let $M$ be a singular semi-Riemannian manifold endowed with a (0,2)(0,2)–tensor field $g$ and the surface of signature change defined as $\mathcal{H}:=\{q\in M\!\!:g\!\!\mid_{q}is\;degenerate\}$. Then $(M,g)$ is a transverse, signature-type changing manifold with a transverse radical if and only if for every $q\in\mathcal{H}$ there exist a neighborhood $U(q)$ and smooth coordinates $(t,x^{1},\ldots,x^{n-1})$ such that $g=-t(dt)^{2}+g_{ij}(t,x^{1},\ldots,x^{n-1})dx^{i}dx^{j}$, for $i,j\in\{1,\ldots,n-1\}$.
@@ 2 Pseudo-causal and pseudo-lightlike curves
In $(M,g)$ be an nn–dimensional manifold on which the metric signature changes from Riemannian, (+,+,,+)($\text{$\cdots$}$,+), to Lorentzian, (,+,,+)($\text{\textminus}$$\text{$\cdots$},+$). In such a setting, the conventional notion of timelike curves ceases to apply globally. From a point located in the Lorentzian region, the hypersurface of signature change may be reached within finite proper time. However, in the Riemannian region no meaningful notion of proper time exists, and hence curves there cannot be classified as timelike, spacelike, or null. Consequently, in manifolds with signature change this gives rise to a novel notion of curves. In order to define those curves we have to make a detour to draw upon the concept of the generalized affine parameter.
@@ 2.1 Properties of the generalized affine parameter
Let $M$ be an nn–dimensional manifold with an affine connection, and let $\gamma\colon J\rightarrow M$ be a $C^{1}$ curve. A smooth vector field $V$ along $\gamma$ is a smooth map $V\colon J\rightarrow TM$ such that $V(t)\in T_{\gamma(t)}M$ for all $t\in J$. Such a vector field $V$ along $\gamma$ is called parallel along $\gamma$ if it satisfies the differential equation
@@ 2.2 Application of the generalized affine parameter in a signature-type changing manifold
Let $M=M_{L}\cup\mathcal{H}\cup M_{R}$ be an nn–dimensional transverse type-changing singular semi-Riemannian manifold with a type-changing metric $g$, and $\mathcal{H}:=\{q\in M\!\!:g\!\!\mid_{q}is\;degenerate\}$ the locus of signature change. We further assume that one component, $M_{L}$, of $M\setminus\mathcal{H}$ is Lorentzian and the other one, $M_{R}$, is Riemannian.
@@
This yields the metric $g=\textrm{sgn}(T)(dT)^{2}+(dx)^{2}$, and for the curve $\gamma$ we have $T=\tan x$. Thus, in the (T,x)(T,x)–coordinates, $\gamma$ is simply the $\tan$–function, with derivative $dT/dx=1/\cos^{2}x$. Consequently, $\gamma$ is timelike in $M_{L}$ , approaching the light cone from timelike infinity and touching it tangentially as $T\rightarrow 0$ (where the derivative becomes $\frac{1}{\cos^{2}(0)}=1$ in the limit). These are precisely the types of curves we aim to exclude in our definition. Note that unlike $(T,x)$, the (t,x)(t,x)–coordinates cover the entire manifold $M$.
@@
Moreover, if the curve $\gamma=(T(s),x(s))$ is parametrized by arc length $s$, then in the (t,x)(t,x)–coordinate system both $\frac{dx}{ds}$ and $\frac{dt}{ds}$ diverge in $M_{L}$, since $1=-(\frac{dT}{ds})^{2}+(\frac{dx}{ds})^{2}$:
@@
While the components of $\gamma^{\prime}$ do not diverge in the (T,x)(T,x)–coordinate system, both $\frac{dx}{ds}$ and $\frac{dt}{ds}$ diverge in $M_{L}$ in the (t,x)(t,x)–coordinates. Because of this dependency of coordinates the criterion of divergence is not useful for defining pseudo-timelike and pseudo-spacelike curves. That is where the coordinate-independent generalized affine parameter comes into play.
@@
Let $M=M_{L}\cup\mathcal{H}\cup M_{R}$ be an nn–dimensional transverse type-changing singular semi-Riemannian manifold, $g$ a type-changing metric, and $\mathcal{H}:=\{q\in M\!\!:g\!\!\mid_{q}is\;degenerate\}$ the locus of signature change. We assume that one component, $M_{L}$, of $M\setminus\mathcal{H}$ is Lorentzian, and the other component, $M_{R}$, is Riemannian.
@@
Revisiting Example 1, we find that both coordinate vector fields, $\frac{\partial}{\partial T}$ and $\frac{\partial}{\partial x}$, are covariantly constant in $M_{L}$ and $M_{R}$ (this is because the Christoffel symbols all vanish in the (T,x)(T,x)–coordinate system). Hence, we can parallel transport $\frac{\partial}{\partial T}$ and $\frac{\partial}{\partial x}$ along any curve in $M_{L}$ and $M_{R}$, with the transport being path-independent (no anholonomy).
@@
reveals that the curve $\gamma$ is not pseudo-timelike as it does not meet the $\varepsilon$–requirement of Definition 8.
@@
Finally, if we modify the previously discussed curve $\gamma$ by keeping the tt–coordinate but stating $x=0$, we get a curve $\alpha$. With the same notation as above, we then get $V^{0}(t)=\sqrt{\mid t\mid}$, $V^{1}(t)=0$ and $\frac{d\mu}{dt}=\sqrt{\mid t\mid}$. Hence, this results in $g(\dot{\hat{\alpha}}(\mu(t)),\dot{\hat{\alpha}}(\mu(t)))=\frac{1}{\mid t\mid}g(\frac{\partial}{\partial t},\frac{\partial}{\partial t})=\frac{t}{\mid t\mid}=-1$ in the Lorentzian region $M_{L}$. The curve $\alpha$ is pseudo-timelike as it obviously does meet the $\varepsilon$–requirement of Definition 8.
@@
The author illustrates the situation in his Figure 1 which contrasts a timelike curve with a timelike almost everywhere curve. The latter one can not be viewed as a timelike curve because it approaches a null vector at its break point. Compared to our setting, however, the culprit here is that the curve is not differentiable at the breaking point. However, if we were to make the curve differentiable by bending its upper section, it still wouldn’t be timelike. Its restriction to the lower region before the inflection point is timelike, but it cannot be extended upwards into a timelike curve. Now, imagine we are not in Minkowski spacetime, but instead, a signature-type change occurs at the (former) inflection point so that the “upper half” of the space becomes Euclidean (in this case, the figure would correspond to the (T,x)(T,x)–coordinates, not the (t,x)(t,x)–coordinates, in the toy model). In this scenario, the curve restricted to the Lorentz sector would be timelike, but after the signature change is “reversed”, it cannot be extended upwards into a timelike curve. In this sense, the entire curve in the signature-changing version of this example is not pseudo-timelike.
@@ 3 Global structure of signature-type changing semi-Riemannian manifolds
[7] A smooth nn–dimensional manifold $M$ is orientable if and only if it has a smooth global nowhere vanishing nn–form (also called a top-ranked form). [^8] An orientation of $M$ is the choice of a continuous pointwise orientation, i.e. the specific choice of a global nowhere vanishing nn–form.
@@
A smooth nn–dimensional manifold $M$ is parallelizable if there exists a set of smooth vector fields $\{V,E_{1},\ldots,E_{n-1}\}$ on $M$, such that at every point $p\in M$ the tangent vectors $\{V(p),E_{1}(p),\ldots,E_{n-1}(p)\}$ provide a basis of the tangent space $T_{p}M$. A specific choice of such a basis of vector fields on $M$ is called an absolute parallelism of $M$.
@@
A signature-type changing manifold $(M,g)$ of dimension $n$ is pseudo-space orientable if and only if it admits a continuous non-vanishing spacelike (n1)(n-1)–frame field on $M_{L}$. This is a set of n1n-1 pointwise orthonormal spacelike vector fields on $M_{L}$. [^13] A pseudo-space orientation of a manifold $(M,g)$ corresponds to the specific choice of a continuous non-vanishing field of orthonormal spacelike (n1)(n-1)–beins on $M_{L}$.
@@
A pseudo-spacetime is a 44–dimensional, pseudo-time oriented, semi-Riemannian manifold with a type-changing metric.
@@
Let $(\mathbb{R}^{n},g)$ be a transverse, signature-type changing nn–manifold with a transverse radical, and let $\mathcal{H}\subset\mathbb{R}^{n}$ be a codimension one closed hypersurface of signature change without boundary. [^15] Here “closed” is meant in the topological sense of “the complement of an open subset of $\mathbb{R}^{n}$” and not in the manifold sense of “a manifold without boundary that is compact.” Then $\mathcal{H}$ is always orientable.
@@
Let $(M,g)$ be a transverse, signature-type changing, oriented, nn–dimensional manifold with a transverse radical, and let $\mathcal{H}\subset M$ be the hypersurface of signature change. Then $\mathcal{H}$ is also oriented.
@@
Let $(M,g)$ be a pseudo-time orientable, transverse, signature-type changing nn–dimensional manifold with a transverse radical. Then in the neighborhood of $\mathcal{H}$ the absolute time function $\mathfrak{h}(t,\mathbf{\hat{x}}):=t$, where $\mathbf{\hat{x}}=(x^{1},\ldots,x^{n-1})$, imposes a natural time direction by postulating that the future corresponds to the increase of the absolute time function. In this way, the time orientation is determined in $M_{L}$.
@@
Respective past-directed curves are defined analogously. Notice that, per assumption, one connected component of $M\setminus\mathcal{H}$ is Riemannian and all other connected components $(M_{L_{\alpha}})_{\alpha\in I}\subseteq M_{L}\subset M$ are Lorentzian. This configuration could (at least locally) potentially allow for a $M_{L}-M_{R}-M_{L}$–sandwich-like structure of $M$, where $\mathcal{H}$ consists of two connected components $(\mathcal{H}_{\alpha})_{\alpha\in\{1,2\}}$. Consequently, this would also imply the existence of two absolute time functions, see Figure 4.
@@
Let $(M,g)$ be a pseudo-time orientable, transverse, signature-type changing nn–dimensional manifold with a transverse radical.

### v3

@@ 1 Introduction
Since a signature-type changing metric is necessarily either degenerate or discontinuous at the locus of signature change [9], we will allow for the metric to become degenerate. Hence, in the present article we will discuss singular semi-Riemannian manifolds for which the metric constitutes a smooth $(0,2)$-tensor field that is degenerate at a subset $\mathcal{H}\subset M$, where the bilinear type of the metric changes upon crossing $\mathcal{H}$.
@@ 1.1 Transverse type-changing singular semi-Riemannian manifold
We align with [25] in requiring that $(M,g)$ be a semi-Riemannian manifold with $\dim M\geq 2$, where $g$ is a smooth, symmetric, degenerate $(0,2)$-tensor on $M$, and $\mathcal{H}:=\{q\in M\!\!:g\!\!\mid_{q}is\;degenerate\}$. This means $\mathcal{H}$ is the locus where the rank of $g$ fails to be maximal. In addition, we assume that one connected component of $M\setminus\mathcal{H}$ is Riemannian, denoted by $M_{R}$, while all other connected components $(M_{L_{\alpha}})_{\alpha\in I}\subseteq M_{L}\subset M$ are Lorentzian, where $M_{L}:=\underset{\alpha\in I}{\bigcup}M_{L_{\alpha}}$represents the Lorentzian domain. Furthermore, we assume throughout that the point set $\mathcal{H}$, where $g$ becomes degenerate is not empty.
@@
    We call the metric $g$ a codimension-$1$ transverse type-changing metric if $d(\det([g_{\mu\nu}]))_{q}\neq 0$ for any $q\in\mathcal{H}$ and any local coordinate system $\xi=(x^{0},\ldots,x^{n-1})$ around $q$. Then we call $(M,g)$ a transverse type-changing singular semi-Riemannian manifold [1, 2, 25]
@@
Let $M$ be a singular semi-Riemannian manifold endowed with a $(0,2)$-tensor field $g$ and the surface of signature change defined as $\mathcal{H}:=\{q\in M\!\!:g\!\!\mid_{q}is\;degenerate\}$. Then $(M,g)$ is a transverse, signature-type changing manifold with a transverse radical if and only if for every $q\in\mathcal{H}$ there exist a neighborhood $U(q)$ and smooth coordinates $(t,x^{1},\ldots,x^{n-1})$ such that $g=-t(dt)^{2}+g_{ij}(t,x^{1},\ldots,x^{n-1})dx^{i}dx^{j}$, for $i,j\in\{1,\ldots,n-1\}$.
@@ 2 Pseudo-causal and pseudo-lightlike curves
In $(M,g)$ be an $n$-dimensional manifold on which the metric signature changes from Riemannian, (+,+,,+)($\text{$\cdots$}$,+), to Lorentzian, (,+,,+)($\text{\textminus}$$\text{$\cdots$},+$). In such a setting, the conventional notion of timelike curves ceases to apply globally. From a point located in the Lorentzian region, the hypersurface of signature change may be reached within finite proper time. However, in the Riemannian region no meaningful notion of proper time exists, and hence curves there cannot be classified as timelike, spacelike, or null. Consequently, in manifolds with signature change this gives rise to a novel notion of curves. In order to define those curves we have to make a detour to draw upon the concept of the generalized affine parameter.
@@ 2.1 Properties of the generalized affine parameter
Let $M$ be an $n$-dimensional manifold with an affine connection, and let $\gamma\colon J\rightarrow M$ be a $C^{1}$ curve. A smooth vector field $V$ along $\gamma$ is a smooth map $V\colon J\rightarrow TM$ such that $V(t)\in T_{\gamma(t)}M$ for all $t\in J$. Such a vector field $V$ along $\gamma$ is called parallel along $\gamma$ if it satisfies the differential equation
@@ 2.2 Application of the generalized affine parameter in a signature-type changing manifold
Let $M=M_{L}\cup\mathcal{H}\cup M_{R}$ be an $n$-dimensional transverse type-changing singular semi-Riemannian manifold with a type-changing metric $g$, and $\mathcal{H}:=\{q\in M\!\!:g\!\!\mid_{q}is\;degenerate\}$ the locus of signature change. We further assume that one component, $M_{L}$, of $M\setminus\mathcal{H}$ is Lorentzian and the other one, $M_{R}$, is Riemannian.
@@
This yields the metric $g=\textrm{sgn}(T)(dT)^{2}+(dx)^{2}$, and for the curve $\gamma$ we have $T=\tan x$. Thus, in the $(T,x)$-coordinates, $\gamma$ is simply the $\tan$-function, with derivative $dT/dx=1/\cos^{2}x$. Consequently, $\gamma$ is timelike in $M_{L}$ , approaching the light cone from timelike infinity and touching it tangentially as $T\rightarrow 0$ (where the derivative becomes $\frac{1}{\cos^{2}(0)}=1$ in the limit). These are precisely the types of curves we aim to exclude in our definition. Note that unlike $(T,x)$, the $(t,x)$-coordinates cover the entire manifold $M$.
@@
Moreover, if the curve $\gamma=(T(s),x(s))$ is parametrized by arc length $s$, then in the $(t,x)$-coordinate system both $\frac{dx}{ds}$ and $\frac{dt}{ds}$ diverge in $M_{L}$, since $1=-(\frac{dT}{ds})^{2}+(\frac{dx}{ds})^{2}$:
@@
While the components of $\gamma^{\prime}$ do not diverge in the $(T,x)$-coordinate system, both $\frac{dx}{ds}$ and $\frac{dt}{ds}$ diverge in $M_{L}$ in the $(t,x)$-coordinates. Because of this dependency of coordinates the criterion of divergence is not useful for defining pseudo-timelike and pseudo-spacelike curves. That is where the coordinate-independent generalized affine parameter comes into play.
@@
Let $M=M_{L}\cup\mathcal{H}\cup M_{R}$ be an $n$-dimensional transverse type-changing singular semi-Riemannian manifold, $g$ a type-changing metric, and $\mathcal{H}:=\{q\in M\!\!:g\!\!\mid_{q}is\;degenerate\}$ the locus of signature change. We assume that one component, $M_{L}$, of $M\setminus\mathcal{H}$ is Lorentzian, and the other component, $M_{R}$, is Riemannian.
@@
Revisiting Example 1, we find that both coordinate vector fields, $\frac{\partial}{\partial T}$ and $\frac{\partial}{\partial x}$, are covariantly constant in $M_{L}$ and $M_{R}$ (this is because the Christoffel symbols all vanish in the $(T,x)$-coordinate system). Hence, we can parallel transport $\frac{\partial}{\partial T}$ and $\frac{\partial}{\partial x}$ along any curve in $M_{L}$ and $M_{R}$, with the transport being path-independent (no anholonomy).
@@
reveals that the curve $\gamma$ is not pseudo-timelike as it does not meet the $\varepsilon$-requirement of Definition 8.
@@
Finally, if we modify the previously discussed curve $\gamma$ by keeping the $t$-coordinate but stating $x=0$, we get a curve $\alpha$. With the same notation as above, we then get $V^{0}(t)=\sqrt{\mid t\mid}$, $V^{1}(t)=0$ and $\frac{d\mu}{dt}=\sqrt{\mid t\mid}$. Hence, this results in $g(\dot{\hat{\alpha}}(\mu(t)),\dot{\hat{\alpha}}(\mu(t)))=\frac{1}{\mid t\mid}g(\frac{\partial}{\partial t},\frac{\partial}{\partial t})=\frac{t}{\mid t\mid}=-1$ in the Lorentzian region $M_{L}$. The curve $\alpha$ is pseudo-timelike as it obviously does meet the $\varepsilon$-requirement of Definition 8.
@@
The author illustrates the situation in his Figure 1 which contrasts a timelike curve with a timelike almost everywhere curve. The latter one can not be viewed as a timelike curve because it approaches a null vector at its break point. Compared to our setting, however, the culprit here is that the curve is not differentiable at the breaking point. However, if we were to make the curve differentiable by bending its upper section, it still wouldn’t be timelike. Its restriction to the lower region before the inflection point is timelike, but it cannot be extended upwards into a timelike curve. Now, imagine we are not in Minkowski spacetime, but instead, a signature-type change occurs at the (former) inflection point so that the “upper half” of the space becomes Euclidean (in this case, the figure would correspond to the $(T,x)$-coordinates, not the $(t,x)$-coordinates, in the toy model). In this scenario, the curve restricted to the Lorentz sector would be timelike, but after the signature change is “reversed”, it cannot be extended upwards into a timelike curve. In this sense, the entire curve in the signature-changing version of this example is not pseudo-timelike.
@@ 3 Global structure of signature-type changing semi-Riemannian manifolds
[7] A smooth $n$-dimensional manifold $M$ is orientable if and only if it has a smooth global nowhere vanishing $n$-form (also called a top-ranked form). [^8] An orientation of $M$ is the choice of a continuous pointwise orientation, i.e. the specific choice of a global nowhere vanishing $n$-form.
@@
A smooth $n$-dimensional manifold $M$ is parallelizable if there exists a set of smooth vector fields $\{V,E_{1},\ldots,E_{n-1}\}$ on $M$, such that at every point $p\in M$ the tangent vectors $\{V(p),E_{1}(p),\ldots,E_{n-1}(p)\}$ provide a basis of the tangent space $T_{p}M$. A specific choice of such a basis of vector fields on $M$ is called an absolute parallelism of $M$.
@@
A signature-type changing manifold $(M,g)$ of dimension $n$ is pseudo-space orientable if and only if it admits a continuous non-vanishing spacelike (n1)(n-1)-frame field on $M_{L}$. This is a set of n1n-1 pointwise orthonormal spacelike vector fields on $M_{L}$. [^13] A pseudo-space orientation of a manifold $(M,g)$ corresponds to the specific choice of a continuous non-vanishing field of orthonormal spacelike (n1)(n-1)-beins on $M_{L}$.
@@
A pseudo-spacetime is a $4$-dimensional, pseudo-time oriented, semi-Riemannian manifold with a type-changing metric.
@@
Let $(\mathbb{R}^{n},g)$ be a transverse, signature-type changing $n$-manifold with a transverse radical, and let $\mathcal{H}\subset\mathbb{R}^{n}$ be a codimension one closed hypersurface of signature change without boundary. [^15] Here “closed” is meant in the topological sense of “the complement of an open subset of $\mathbb{R}^{n}$” and not in the manifold sense of “a manifold without boundary that is compact.” Then $\mathcal{H}$ is always orientable.
@@
Let $(M,g)$ be a transverse, signature-type changing, oriented, $n$-dimensional manifold with a transverse radical, and let $\mathcal{H}\subset M$ be the hypersurface of signature change. Then $\mathcal{H}$ is also oriented.
@@
Let $(M,g)$ be a pseudo-time orientable, transverse, signature-type changing $n$-dimensional manifold with a transverse radical. Then in the neighborhood of $\mathcal{H}$ the absolute time function $\mathfrak{h}(t,\mathbf{\hat{x}}):=t$, where $\mathbf{\hat{x}}=(x^{1},\ldots,x^{n-1})$, imposes a natural time direction by postulating that the future corresponds to the increase of the absolute time function. In this way, the time orientation is determined in $M_{L}$.
@@
Respective past-directed curves are defined analogously. Notice that, per assumption, one connected component of $M\setminus\mathcal{H}$ is Riemannian and all other connected components $(M_{L_{\alpha}})_{\alpha\in I}\subseteq M_{L}\subset M$ are Lorentzian. This configuration could (at least locally) potentially allow for a $M_{L}-M_{R}-M_{L}$-sandwich-like structure of $M$, where $\mathcal{H}$ consists of two connected components $(\mathcal{H}_{\alpha})_{\alpha\in\{1,2\}}$. Consequently, this would also imply the existence of two absolute time functions, see Figure 4.
@@
Let $(M,g)$ be a pseudo-time orientable, transverse, signature-type changing $n$-dimensional manifold with a transverse radical.